OpenStudy (anonymous):

How do I find the product of (4p-1)^2

OpenStudy (amistre64):

Finding the product means that you multiply 2 or more things together. Your example is the same as trying to find (8)^2. This means that we multiply (8) by itself 2 times. (8)*(8) = 64. Lets try your problem (4p-1)^2 means: (4p-1)(4p-1) = (4p-1)(4p-1) (4p-1)(4p) + (4p-1)(-1) = (4p-1)(4p-1) (4p)(4p) + (4p)(-1) + (4p)(-1) + (-1)(-1) = (4p-1)(4p-1) (16p^2) + (-4p) + (-4p) + (1) = (4p-1)(4p-1) (16p^2) + (-8p) + (1) = (4p-1)(4p-1) 16p^2 -8p +1 = (4p-1)(4p-1)

OpenStudy (amistre64):

The short cut is: (a+b)^2 = (a^2 +2ab +b^2) another short cut is FOIL (Firsts + Outsides + Insides + Lasts) (a+b)(a+b) = a^2 + ab + ab + b^2 = a^2 +2ab +b^2 but that is how you would find it if you forgot the short cuts

OpenStudy (anonymous):

using formula, (a-b)^2=a^2-2ab+b^2 a=4p,b=1 =16p^2-8p+1